# Deep Multimodal Learning with Missing Modality: A Survey

### v3

@@ 3.1.2. Modality Generation Methods
Individual Modality Generation Methods train an individual generative model for each modality in case any modality is missing, as shown in Figure 5(a). Early works used Gaussian processes (Mario Christoudias et al., 2010) or Boltzmann machines (Sohn et al., 2014) to generate missing modalities from available data. With deep learning, models like Auto-Encoders (AEs) (Hinton and Zemel, 1993) can be used for missing modality generation. Li et al. (Li et al., 2014) used 3D-CNN to generate positron emission tomography (PET) data from magnetic resonance imaging (MRI) inputs. Chen et al. (Chen et al., 2024b) addressed missing modalities in MRI segmentation by training U-Net models to generate other two modalities from MRI. A recent work (Ma et al., 2021b) used AEs as one of the baselines to complete datasets by training one AE per modality. In domain adaptation, Zhang et al. (Zhang et al., 2021) proposed a U-Net-based Multi-Modality Data Generation module with domain adversarial learning to generate each missing modality by learning domain-invariant features.
@@
GANs significantly improved image generation quality by using a generator to create realistic data and a discriminator to distinguish it from real data. Researchers therefore began replacing above models with GANs for modality generation. For example, GANs generate missing modalities using latent representations of existing ones in breast cancer prediction (Arya and Saha, 2021). In remote sensing, Bischke et al. (Bischke et al., 2018) used GANs to generate depth data, improving segmentation over RGB-only models. GANs were also used in robotic object recognition models to help generate missing RGB and depth images(Gunasekar et al., 2020). Recent studies (Ma et al., 2021b) show that GANs outperform AEs in generating more realistic missing modalities and can lead to better downstream-task model performance. Recently, the introduction of Diffusion models has further improved image generation quality. Wang et al. proposed the IMDer method (Wang et al., 2024a), which uses available modalities as conditions to make diffusion models generate missing modalities. Experiments showed it can reduce semantic ambiguity between recovered and missing modalities and achieves good generalization performance than previous works.
@@
Unified Original Data Generative Methods train a unified model that can generate all modalities simultaneously (Figure 5(b)). One representative model is Cascade AE (Tran et al., 2017), which stacks AEs to capture the differences between missing and existing modalities for generating all missing modalities. Recently, Zhang et al. (Zhang et al., 2023b), have attempted to use attention and max-pooling to integrate features of existing modalities, enabling modality-specific decoders to accept available modalities and generate other missing modalities. Experiments demonstrated that this method is more effective than using max-pooling alone (Chartsias et al., 2017) to integrate features from any number of modalities for generating all missing modalities together. Above methods for generating missing modalities can mitigate performance degradation to some extent. However, when facing a scenraio where one of the modalities is severely missing, training a generator that can produce high-quality missing modalities remains challenging. Additionally, the modality generation model significantly increases storage and computational requirements. As the number of modalities grows, the complexity of these generative models also increases, further complicating the training process and resource demands.
@@
Figure 6. Illustration of coordinated-representation methods, which impose constraints to make the learned representations semantically consistent. Assuming that modality 2 is missing, these methods use constraints terms between the features of different modalities.
3.2. Representation-Focused Models
@@
3.2.1. Coordinated-Representation Methods
@@
Some methods train models by coordinating the correlation between different modal features. For example, Wang et al. (Wang et al., 2020a) use a Deep Canonical Correlation Analysis (CCA) module to maximize feature associations of available modalities by using Canonical Correlation Coefficient, which enables training on incomplete datasets. Ma et al. (Ma et al., 2021c) proposed a Maximum likelihood function to characterize conditional distributions of full-modality and missing-modality samples during training. Other research efforts (Liu et al., 2021a) have added constraints based on the Hilbert-Schmidt Independence Criterion (HSIC), which helps models learn how to complete missing modality features by enforcing independence between irrelevant features while maximizing the dependence between relevant ones. These methods (Matsuura et al., 2018; Ma et al., 2021a; Li et al., 2022a) aid models in learning how to complete missing modality features or training on incomplete datasets by learning the similarity or correlation between features. A drawback of above methods is that they perform well only when two or three modalities are used as input (Zhao et al., 2024). Even when trained on a dataset with missing modalities, some methods still struggle to effectively address the missing modality problem during testing.
@@
Figure 7. The general idea of the arithmetic operation-based representation composition methods. The representation of any number of modalities can be combined to a fixed dimension that is the same as the dimension required by the subsequent DNN layer.
3.2.2. Representation Composition Methods
@@
There are two types of representation composition methods, which are explained below. Retrieval-Based Representation Composition Methods attempt to recover the missing modality representation by retrieving the modality data from existing samples, similarly to the modality composition method in Section 3.1.1. They typically use pre-trained feature extractors to generate features from available samples, storing them in a feature pool (Zhi et al., 2024; Sun et al., 2024a). Cosine similarity is then used to retrieve matched features for the input sample, with the average of these missing-modality features in top-K samples filling the representation of missing modalities. Additionally, some methods, such as Missing Modality Feature Generation (Wang et al., 2023c), replace missing modality features by averaging the representations of available modalities, assuming similar feature distributions across modalities.
@@
Arithmetic Operation-Based Representation Composition Methods can flexibly fuse any number of modality representations through arithmetic operations such as pooling without learnable parameters (Figure 7). Researchers (Zhou et al., 2021; Wang et al., 2023c) have fused features through operations like average/max pooling or addition, offering low computational complexity and efficiency. However, drawbacks include potential loss of important information. To address this, merge operations (Delgado-Escano et al., 2021) use a sign-max function, which selects the highest absolute value from feature vectors, yielding better results by preserving both positive and negative activation values. TMFormer (Zhang et al., 2024) introduces token merging based on cosine similarity. Similar approaches (Zheng et al., 2024) focus on selecting key vectors or merging tokens to handle missing modalities.
@@
Figure 8. Description of two typical representation generation methods. We assume modality 2 is missing here. (a) Indirect-to-task representation generation methods are supervised by the loss of two tasks, namely: auxiliary reconstruction and downstream tasks. Note that these two tasks sometimes can be trained separately, first training the reconstruction task, discarding the generator, and then training the downstream task, such as the training progress of ActionMAE (Woo et al., 2023). While (b) direct-to-task representation generation methods directly use a modality generation network (“GEN-2” in figure) to generate representations of the missing modality 2 during training and inference without reconstructing the modality 2.
3.2.3. Representation Generation Methods
@@
Before Fusion methods receive the features from modality-specific encoders directly into each specific reconstruction decoder for missing modality reconstruction. For example, MGP-VAE (Hamghalam et al., 2021) uses output of the VAE encoder of each modality as the input of each reconstruction decoder and the fused output for Gaussian Process prediction. Similarly, PFNet (Zheng et al., 2021) employs RGB and thermal-infrared modalities for pedestrian re-identification, feeding outputs of each encoder into corresponding decoders before fusing them together. Li et al. (Li et al., 2024b) adopted a similar approach for brain tumor segmentation, reconstructing each modality before fusing encoder outputs.
After Fusion methods use the fused features from modality-specific encoders as the input of reconstruction decoder. Tsai et al. (Tsai et al., 2018) introduced a Multimodal Factorization Model where modality-specific encoder-decoders reconstruct the missing modality, while the downstream task module employs a separate multimodal encoder-decoder. To recover the missing modality features, the outputs of the downstream task model’s encoder are not only used for predictions of the downstream task decoder, but also fused with the outputs of each modality-specific encoder for inputting to the each corresponding reconstruction decoder. Chen et al. (Chen et al., 2019) also decoupled the encoders for reconstruction and downstream tasks, showing that disentangling features helped eliminate irrelevant features. Jeong et al. (Jeong et al., 2022) proposed a simpler approach using multi-level skip connections in a U-Net architecture to fuse features for missing modality reconstruction, achieving better generalization in brain tumor segmentation than Chen et al. (Chen et al., 2019). Further, Zhou et al. (Zhou, 2023) and Sun et al. (Sun et al., 2024b) demonstrated effective cross-modal attention fusion techniques could help enhance reconstruction and segmentation. Most of these methods train reconstruction and downstream tasks concurrently. However, some approaches, such as ActionMAE (Woo et al., 2023) and M3AE (Liu et al., 2023c), first train reconstruction tasks and then fine-tune the pre-trained encoders for downstream tasks. Although these methods achieve strong performance, they still require the missing modality data to be present during training to supervise the reconstruction decoders.
@@
In tasks like sentiment analysis, a translation module can map available-modality representations to the missing ones, with L2 loss applied to align outputs with those from networks trained on missing modalities (Huan et al., 2023). A notable recent work, SMIL (Ma et al., 2021b), employs Bayesian learning and meta-regularization to directly generate missing modality representations. SMIL has shown strong generalization on datasets with a high rate of missing modality samples. Subsequent work (Li et al., 2022d) addressed challenges such as distribution inconsistency, failure to capture specific modality information, and lack of correspondence between generated and existing modalities. Researchers have also explored cross-modal distribution transformation methods to align distributions of generated and existing modality representations, enhancing discriminative ability (Wang et al., 2023b). In audio-visual question answering (AVQA), Park et al. (Park et al., 2024) proposed a Relation-aware Missing Modal Generator (RMM) that consist of to generate pseudo features of missing modalities, improving robustness and accuracy. The RMM consists of a visual generator, auditory generator, and text generator, which generate missing auditory representations by analyzing the correlations between visual and textual modalities and utilizing learnable parameter vectors (slots) for reconstruction and synthesis of missing-modality features.
Recently, to address the potential absence of different modalities in sentiment analysis, Guo et al. (Guo et al., 2024) proposed a Missing Modality Generation Module for prompt learning. This module maps prompts of available modalities to prompts of missing modalities through a set of projection layers. Lin et al. (Lin et al., 2024) proposed Uncertainty Estimation Module which can identify useless tokens in different modality tokens to facilitate better use of U-Adapter-assisted pre-trained models to better utilize the tokens of available modalities when downstream tasks face missing modalities.
@@
In summary, representation generation methods are limited by the imbalance of missing modalities in the training dataset or the limited number of full-modality samples, which may lead to training procedures that overfit to the existing modalities. In addition, indirect-to-task representation generation methods can access the missing modality data of the missing modality samples during training, which means that the training dataset is actually complete, so this type of method usually performs better than direct-to-task representation generation methods.
4. Methodologies in Strategy Design Aspect
4.1. Architecture-Focused Models
@@
Figure 9. Two typical attention fusion methods, assuming modality 2 as the missing modality. (a) The intra-modality attention (“Intra-Atten” in figure) of missing modality 2 can be skipped. Since each Intra-Atten of different modalities shares the fusion representation (like learnable query in (Ge et al., 2023) or bottleneck tokens in (Nagrani et al., 2021)), after calculating them one by one, the output representation can be regarded as the fusion of available modalities. (b) “Missing Modality Mask” is the custom mask that is generated according to the missing modality 2 and can help inter-modality attention (“Inter-Atten” in figure) ignore the missing modality 2. By setting the tokens of modality 2 to zero or negative infinity through masking, we can force the attention mechanism to ignore the missing modality 2.
4.1.1. Attention Based Methods
@@
Attention Fusion Methods have the powerful ability to capture key features and can be seen as plug-and-play modules. We categorize them into two types: intra- and inter-modality attention methods. Intra-Modality Attention Methods compute attention for each modality independently before fusing them, as shown in Figure 9(a). This approach focuses on relationships within a single modality, and the fusion between modalities is achieved by sharing partial information. For instance, in 3D detection, the BEV-Evolving Decoder (Ge et al., 2023) handles sensor failures by sharing same BEV-query with each modality-specific attention modules, allowing fusion of any number of modalities. Similarly, in clinical diagnosis, Lee et al. (Lee et al., 2023a) proposed modality-aware attention to perform intra-modality attention and predict decisions by using shared bottleneck fusion tokens (Nagrani et al., 2021). Inter-Modality Attention Methods, often based on masked attention, treat missing modality features as masked vectors (using zero or negative infinity values) to better capture dependencies across available modalities, as illustrated in Figure 9(b). Unlike conventional cross-modal attention, masked attention models share the same parameters across all embeddings, allowing flexible handling of missing modalities. For example, Qian et al. (Qian and Wang, 2023) designed an attention mask matrix to ignore missing modalities, improving model robustness. Similarly, DrFuse(Yao et al., 2024) decouples modalities into specific and shared representations, using the shared ones to replace missing modalities, with a custom mask matrix to help model ignore the specific representations of the missing modality.
@@
Figure 10. Description of general idea of joint representation learning methods, assuming Modality 2 is the missing modality. The missing modality 2 tokens can be replaced by specific masked tokens. The modality encoders (red dashed box) can be replaced by linear projection layers in some methods.
@@
Transformer Based Methods can be divided into two types: joint representation learning (JRL) and parameter efficient learning (PEL) according to full parameter training and a small amount of parameter fine-tuning. Since transformers have long context lengths to handle many feature tokens, the multimodal transformer can learn joint representations from any number of modality tokens (Figure 10). Some methods first use modality encoders to extract feature tokens from available modalities and then feed them into a multimodal transformer. Gong et al. (Gong et al., 2023) introduced an egocentric multimodal task, proposing a transformer-based fusion module with a flexible number of modality tokens and a cross-modal contrast alignment loss to map features into a common space. Similarly, Mordacq et al. (Mordacq et al., 2024) leveraged Masked Multimodal Transformers, treating missing modalities as masked tokens for robust JRL. Ma et al. (Ma et al., 2022) proposed an optimal fusion strategy search strategy within the multimodal transformer to help find the best fusion method for different missing/full-modality representations. Radosavovic et al. (Radosavovic et al., 2024) introduced specially designed mask tokens for an autoregressive transformer to manage missing modalities, successfully deploying the model in real-world scenarios.
@@
With the rise of pre-trained transformer models, PEL methods have been developed to fine-tune these models by training few parameters. Two common PEL methods for pre-trained models are prompt and adapter tuning. Initially used in natural language processing, prompt tuning optimizes input prompts while keeping model parameters fixed. It has since been extended to multimodal models. Jang et al. (Jang et al., 2024) introduced modality-specific prompts to address limitations in earlier methods, which merge when all modalities are present and allow the model to update all learnable prompts during training. Liu et al. (Liu et al., 2024b) further improved this by proposing Fourier Prompt, which uses fast Fourier transform to encode global spectral information of available modalities into learnable prompt tokens which can be used to supplement missing-modality features, enabling cross-attention with feature tokens to address missing modalities. Adapter tuning, on the other hand, involves inserting lightweight adapter layers (e.g., MLPs) into pre-trained models to adapt to new tasks without modifying the original parameters. Qiu et al. (Qiu et al., 2023) proposed a method that used a classifier to identify different missing-modality cases and used intermediate features of that classifier as the missing-modality prompt to cooperate with lightweight Adapters to address missing modality problems.
@@
Figure 11. Description of representation distillation methods. Here we set modality 2 is missing. In order to distinguish the representations of teachers and students, we add “-T” and “-S” after their representations in the figure. We refer to intermediate distillation as using features from any combination of intermediate layers within models.
@@
Figure 12. Description of two process-based distillation methods. (a) “EMA” means exponential moving average. To demonstrate the general architecture of Mean Teacher Distillation, we set the loss on the logits representation. (b) Self-distillation methods for missing modality problems involve the use of models that act as teachers and students, where these models use their own soft labels/representations from each branch to refine themselves during training.
4.1.2. Distillation Based Methods

### v4

@@ 3.1.2 Modality Generation Methods
Individual Modality Generation Methods train an individual generative model for each modality in case any modality is missing, as shown in figure 5(a). Early works used Gaussian processes (Mario Christoudias et al., 2010) or Boltzmann machines (Sohn et al., 2014) to generate missing modalities from available data. With deep learning, models like Auto-Encoders (AEs) (Hinton and Zemel, 1993) can be used for missing modality generation. Li et al. (2014) used 3D-CNN to generate positron emission tomography (PET) data from magnetic resonance imaging (MRI) inputs. Chen et al. (2024a) addressed missing modalities in MRI segmentation by training U-Net models to generate other two modalities from MRI. A recent work (Ma et al., 2021c) used AEs as one of the baselines to complete datasets by training one AE per modality. In domain adaptation, Zhang et al. (2021) proposed a U-Net-based Multi-Modality Data Generation module with domain adversarial learning to generate each missing modality by learning domain-invariant features. In immunohistochemical, DeepLIIF (Ghahremani et al., 2022) uses ResNet to generate the various modalities such as Hematoxylin, mpIF DAPI, mpIF Lap2, and mpIF Ki67. HE2RNA (Schmauch et al., 2020) is a cross-modal model that predicts gene expression (transcriptomic modality) from H&E whole-slide images (visual modality), generates spatialized gene-expression heatmaps, and transfers this representation to downstream clinical tasks.
@@
GANs significantly improved image generation quality by using a generator to create realistic data and a discriminator to distinguish it from real data. Researchers therefore began replacing above models with GANs for modality generation. For example, GANs generate missing modalities using latent representations of existing ones in breast cancer prediction (Arya and Saha, 2021). In remote sensing, Bischke et al. used GANs to generate depth data, improving segmentation over RGB-only models (Bischke et al., 2018). GANs were also used in robotic object recognition models to help generate missing RGB and depth images (Gunasekar et al., 2020). Recent studies (Ma et al., 2021c) show that GANs outperform AEs in generating more realistic missing modalities and can lead to better downstream-task model performance. SensorGAN (Hussein and Bhat, 2024) introduces a GAN model that restores absent sensor channels, directly targeting missing-modality recovery in wearable human-activity-recognition pipelines. Recently, the introduction of Diffusion models has further improved image generation quality. Wang et al. proposed the IMDer method (Wang et al., 2024c), which uses available modalities as conditions to make diffusion models generate missing modalities. Experiments showed it can reduce semantic ambiguity between recovered and missing modalities and achieves good generalization performance than previous works.
@@
Unified Original Data Generative Methods train a unified model that can generate all modalities simultaneously (figure 5(b)). One representative model is Cascade AE (Tran et al., 2017), which stacks AEs to capture the differences between missing and existing modalities for generating all missing modalities. Recently, Zhang et al. (2023c), have attempted to use attention and max-pooling to integrate features of existing modalities, enabling modality-specific decoders to accept available modalities and generate other missing modalities. Experiments demonstrated that this method is more effective than using max-pooling alone (Chartsias et al., 2017) to integrate features from any number of modalities for generating all missing modalities together. Above methods for generating missing modalities can mitigate performance degradation to some extent. However, when facing a scenraio where one of the modalities is severely missing, training a generator that can produce high-quality missing modalities remains challenging. Additionally, the modality generation model significantly increases storage and computational requirements. As the number of modalities grows, the complexity of these generative models also increases, further complicating the training process and resource demands.
@@
Figure 6: Illustration of coordinated-representation methods, which impose constraints to make the learned representations semantically consistent. Assuming that modality 2 is missing, these methods use constraints terms between the features of different modalities.
3.2 Representation-Focused Models
@@
3.2.1 Coordinated-Representation Methods
@@
Some methods train models by coordinating the correlation between different modal features. For example, Wang et al. (2020b) use a Deep Canonical Correlation Analysis (CCA) module to maximize feature associations of available modalities by using Canonical Correlation Coefficient, which enables training on incomplete datasets. Ma et al. (2021b) proposed a Maximum likelihood function to characterize conditional distributions of full-modality and missing-modality samples during training. Other research efforts (Liu et al., 2021b) have added constraints based on the Hilbert-Schmidt Independence Criterion (HSIC), which helps models learn how to complete missing modality features by enforcing independence between irrelevant features while maximizing the dependence between relevant ones. These methods (Matsuura et al., 2018; Ma et al., 2021a; Li et al., 2022a) aid models in learning how to complete missing modality features or training on incomplete datasets by learning the similarity or correlation between features. A drawback of the above methods is that they perform well only when two or three modalities are used as input (Zhao et al., 2024), and their effectiveness tends to degrade as the number of modalities increases. Even when trained on a dataset with missing modalities, some methods still struggle to effectively address unseen missing modality combinations or cases during testing.
@@
Figure 7: The general idea of the arithmetic operation-based representation composition methods. The representation of any number of modalities can be combined to a fixed dimension that is the same as the dimension required by the subsequent DNN layer.
3.2.2 Representation Composition Methods
@@
There are two types of representation composition methods, which are explained below. Retrieval-Based Representation Composition Methods attempt to recover the missing modality representation by retrieving the modality data from existing samples, similarly to the modality composition method in section 3.1.1. They typically use pre-trained feature extractors to generate features from available samples, storing them in a feature pool (Zhi et al., 2024; Sun et al., 2024b). Cosine similarity is then used to retrieve matched features for the input sample, with the average of these missing-modality features in top-K samples filling the representation of missing modalities. Additionally, some methods, such as Missing Modality Feature Generation (Wang et al., 2023b), replace missing modality features by averaging the representations of available modalities, assuming similar feature distributions across modalities.
@@
Arithmetic Operation-Based Representation Composition Methods can flexibly fuse any number of modality representations through arithmetic operations such as pooling without learnable parameters (figure 7). Researchers (Zhou et al., 2021; Wang et al., 2023b) have fused features through operations like average/max pooling or addition, offering low computational complexity and efficiency. However, drawbacks include potential loss of important information. To address this, merge operations (Delgado-Escano et al., 2021) use a sign-max function, which selects the highest absolute value from feature vectors, yielding better results by preserving both positive and negative activation values. TMFormer (Zhang et al., 2024) introduces token merging based on cosine similarity. Similar approaches (Zheng et al., 2024) focus on selecting key vectors or merging tokens to handle missing modalities. A recent work (Sarkar et al., 2025) propose to perform weighted fusion of these pooled features to learn a fixed-size multimodal embedding
@@
Figure 8: Description of two typical representation generation methods. We assume modality 2 is missing here. (a) Indirect-to-task representation generation methods are supervised by the loss of two tasks, namely: auxiliary reconstruction and downstream tasks. Note that these two tasks sometimes can be trained separately, first training the reconstruction task, discarding the generator, and then training the downstream task, such as the training progress of ActionMAE (Woo et al., 2023). While (b) direct-to-task representation generation methods directly use a modality generation network (“GEN-2” in figure) to generate representations of the missing modality 2 during training and inference without reconstructing the modality 2.
3.2.3 Representation Generation Methods
@@
Before Fusion methods receive the features from modality-specific encoders directly into each specific reconstruction decoder for missing modality reconstruction. For example, MGP-VAE (Hamghalam et al., 2021) uses output of the VAE encoder of each modality as the input of each reconstruction decoder and the fused output for Gaussian Process prediction. Similarly, PFNet (Zheng et al., 2021) employs RGB and thermal-infrared modalities for pedestrian re-identification, feeding outputs of each encoder into corresponding decoders before fusing them together. Li et al. (2024b) adopted a similar approach for brain tumor segmentation, reconstructing each modality before fusing encoder outputs.
After Fusion methods use the fused features from modality-specific encoders as the input of reconstruction decoder. Tsai et al. (2018) introduced a Multimodal Factorization Model where modality-specific encoder-decoders reconstruct the missing modality, while the downstream task module employs a separate multimodal encoder-decoder. To recover the missing modality features, the outputs of the downstream task model’s encoder are not only used for predictions of the downstream task decoder, but also fused with the outputs of each modality-specific encoder for inputting to the each corresponding reconstruction decoder. Chen et al. (2019) and Li et al. (2025a) also decoupled the encoders for reconstruction and downstream tasks, showing that disentangling features helped eliminate irrelevant features. Jeong et al. (2022) proposed a simpler approach using multi-level skip connections in a U-Net architecture to fuse features for missing modality reconstruction, achieving better generalization in brain tumor segmentation than (Chen et al., 2019). Further, Zhou (2023) and Sun et al. (2024a) demonstrated effective cross-modal attention fusion techniques could help enhance reconstruction and segmentation. Most of these methods train reconstruction and downstream tasks concurrently. However, some approaches, such as ActionMAE (Woo et al., 2023) and M3AE (Liu et al., 2023a), first train reconstruction tasks and then fine-tune the pre-trained encoders for downstream tasks.
Although these methods achieve strong performance, they still require the missing modality data to be present during training to supervise the reconstruction decoders.
@@
In tasks like sentiment analysis, a translation module can map available-modality representations to the missing ones, with L2 loss applied to align outputs with those from networks trained on missing modalities (Huan et al., 2023). A notable recent work, SMIL (Ma et al., 2021c), employs Bayesian learning and meta-regularization to directly generate missing modality representations. SMIL has shown strong generalization on datasets with a high rate of missing modality samples. Subsequent work (Li et al., 2022d) addressed challenges such as distribution inconsistency, failure to capture specific modality information, and lack of correspondence between generated and existing modalities. Researchers have also explored cross-modal distribution transformation methods to align distributions of generated and existing modality representations, enhancing discriminative ability (Wang et al., 2023d). In audio-visual question answering (AVQA), Park et al. (2024) proposed a Relation-aware Missing Modal Generator (RMM) that consist of to generate pseudo features of missing modalities, improving robustness and accuracy. The RMM consists of a visual generator, auditory generator, and text generator, which generate missing auditory representations by analyzing the correlations between visual and textual modalities and utilizing learnable parameter vectors (slots) for reconstruction and synthesis of missing-modality features.
Recently, to address the potential absence of different modalities in sentiment analysis, Guo et al. (2024) proposed a Missing Modality Generation Module for prompt learning. This module maps prompts of available modalities to prompts of missing modalities through a set of projection layers. Lin et al. (2024) proposed Uncertainty Estimation Module which can identify useless tokens in different modality tokens to facilitate better use of U-Adapter-assisted pre-trained models to better utilize the tokens of available modalities when downstream tasks face missing modalities. Liao et al. (2025) explore a variety of different conditions of sensor impairment are and the capabilities of various existing multimodal segmentation models are evaluated. Park et al. (2025) use the Multi-modal Mixture of Expert, which not only enables the model to dynamically handle multi-modality, but also allows different branches to focus on their own modalities by sharing a query.
@@
In summary, representation generation methods are limited by the imbalance of missing modalities in the training dataset or the limited number of full-modality samples, which may lead to training procedures that overfit to the existing modalities. In addition, indirect-to-task representation generation methods can access complete multimodal data samples during training, which means they can simulate arbitrary missing modality cases during training by dropping modalities. As a result, they generally achieve better performance than direct-to-task methods, which must handle incomplete multimodal data samples during training.
4 Methodologies in Strategy Design Aspect
4.1 Architecture-Focused Models
@@
Figure 9: Two typical attention fusion methods, assuming modality 2 as the missing modality. (a) The intra-modality attention (“Intra-Atten” in figure) of missing modality 2 can be skipped. Since each Intra-Atten of different modalities shares the fusion representation (like learnable query in (Ge et al., 2023) or bottleneck tokens in (Nagrani et al., 2021)), after calculating them one by one, the output representation can be regarded as the fusion of available modalities. (b) “Missing Modality Mask” is the custom mask that is generated according to the missing modality 2 and can help inter-modality attention (“Inter-Atten” in figure) ignore the missing modality 2. By setting the tokens of modality 2 to zero or negative infinity through masking, we can force the attention mechanism to ignore the missing modality 2.
4.1.1 Attention Based Methods
@@
Attention Fusion Methods have the powerful ability to capture key features and can be seen as plug-and-play modules. We categorize them into two types: intra- and inter-modality attention methods. Intra-Modality Attention Methods compute attention for each modality independently before fusing them, as shown in figure 9(a). This approach focuses on relationships within a single modality, and the fusion between modalities is achieved by sharing partial information. For instance, in 3D detection, the BEV-Evolving Decoder (Ge et al., 2023) handles sensor failures by sharing same BEV-query with each modality-specific attention modules, allowing fusion of any number of modalities. Similarly, in clinical diagnosis, Lee et al. (2023a) proposed modality-aware attention to perform intra-modality attention and predict decisions by using shared bottleneck fusion tokens (Nagrani et al., 2021). Inter-Modality Attention Methods, often based on masked attention, treat missing modality features as masked vectors (using zero or negative infinity values) to better capture dependencies across available modalities, as illustrated in figure 9(b). Unlike conventional cross-modal attention, masked attention models share the same parameters across all embeddings, allowing flexible handling of missing modalities. For example, Qian and Wang (2023) designed an attention mask matrix to ignore missing modalities, improving model robustness. Similarly, DrFuse (Yao et al., 2024) decouples modalities into specific and shared representations, using the shared ones to replace missing modalities, with a custom mask matrix to help model ignore the specific representations of the missing modality.
@@
Figure 10: Description of general idea of joint representation learning methods, assuming Modality 2 is the missing modality. The missing modality 2 tokens can be replaced by specific masked tokens. The modality encoders (red dashed box) can be replaced by linear projection layers in some methods.
@@
Transformer Based Methods can be divided into two types: joint representation learning (JRL) and parameter efficient learning (PEL) according to full parameter training and a small amount of parameter fine-tuning. Since transformers have long context lengths to handle many feature tokens, the multimodal transformer can learn joint representations from any number of modality tokens (figure 10). Some methods first use modality encoders to extract feature tokens from available modalities and then feed them into a multimodal transformer. Gong et al. (2023) introduced an egocentric multimodal task, proposing a transformer-based fusion module with a flexible number of modality tokens and a cross-modal contrast alignment loss to map features into a common space. Similarly, Mordacq et al. (2024) leveraged Masked Multimodal Transformers, treating missing modalities as masked tokens for robust JRL. Ma et al. (2022) proposed an optimal fusion strategy search strategy within the multimodal transformer to help find the best fusion method for different missing/full-modality representations. Radosavovic et al. (2024) introduced specially designed mask tokens for an autoregressive transformer to manage missing modalities, successfully deploying the model in real-world scenarios.
@@
With the rise of pre-trained transformer models, PEL methods have been developed to fine-tune these models by training few parameters. Two common PEL methods for pre-trained models are prompt and adapter tuning. Initially used in natural language processing, prompt tuning optimizes input prompts while keeping model parameters fixed. It has since been extended to multimodal models. Jang et al. (2024) introduced modality-specific prompts to address limitations in earlier methods, which merge when all modalities are present and allow the model to update all learnable prompts during training. Liu et al. (2024b) further improved this by proposing Fourier Prompt, which uses fast Fourier transform to encode global spectral information of available modalities into learnable prompt tokens which can be used to supplement missing-modality features, enabling cross-attention with feature tokens to address missing modalities. Adapter tuning, on the other hand, involves inserting lightweight adapter layers (e.g., MLPs) into pre-trained models to adapt to new tasks without modifying the original parameters. Qiu et al. (2023) proposed a method that used a classifier to identify different missing-modality cases and used intermediate features of that classifier as the missing-modality prompt to cooperate with lightweight Adapters to address missing modality problems.
@@
Figure 11: Description of representation distillation methods. Here we set modality 2 is missing. In order to distinguish the representations of teachers and students, we add “-T” and “-S” after their representations in the figure. We refer to intermediate distillation as using features from any combination of intermediate layers within models.
@@
Figure 12: Description of two process-based distillation methods. (a) “EMA” means exponential moving average. To demonstrate the general architecture of Mean Teacher Distillation, we set the loss on the logits representation. (b) Self-distillation methods for missing modality problems involve the use of models that act as teachers and students, where these models use their own soft labels/representations from each branch to refine themselves during training.
4.1.2 Distillation Based Methods